OpenStudy (anonymous):

6 ∫(x+4)√(6-2x) anyone? 2

OpenStudy (zarkon):

u=6-2x ...

OpenStudy (zarkon):

your limits are not good for this problem

OpenStudy (anonymous):

i got that far... and du=-2dx... but how does that work into the eqn

OpenStudy (zarkon):

are you sure your typed the problem correctly?

OpenStudy (anonymous):

should have dx after √(6-2x)

OpenStudy (zarkon):

\[\sqrt{6-2x}\] is a complex when x>3

OpenStudy (zarkon):

* a complex number

OpenStudy (zarkon):

are you sure that the limits of integration are from x=2 to x=6?

OpenStudy (anonymous):

as written on my final review sheet

OpenStudy (zarkon):

then it is a typo

OpenStudy (anonymous):

not surprising, but could I integrate (x+4) as x^2/2+4x and √(6-2x) as (6-2x)^1/2...[(6-2x)^3/2]/3/2

OpenStudy (zarkon):

if you just wanted to compute \[\int(x+4)\sqrt{6-2x}dx\] u=6-2x du=-2dx \[x=\frac{6-u}{2}\] \[\int(\frac{6-u}{2}+4)\sqrt{u}\frac{du}{-2}\] expand...then integrate term by term

OpenStudy (anonymous):

it doesnt work as a definite integral to find area under curve of the eqn. on the interval (2,6)

OpenStudy (zarkon):

no

OpenStudy (anonymous):

ok thanks, a typo wouldn't be unusual